The 2018 Income Statement and Balance Sheet of Tampa Clinic, a not-for-profit organization, are presented below.

Tampa Clinic

Income Statement

Year Ended December 31, 2018

Operating Revenues:

Patient service revenue                                    $524,630

Less provision for bad debts                              (27,630)

   Net patient service revenue                           $497,000

Other revenue                                                     10,000

Net operating revenues                                    $507,000

Expenses:

Salaries and benefits                                         $231,950

Supplies                                                               76,050

Insurance                                                             16,700

Depreciation                                                          8,000

Interest                                                                41,000

Total expenses                                                 $373,700

Operating income                                            $133,300

Nonoperating income:

Investment income                                          $    2,700

Net income                                                      $136,000

Tampa Clinic

Balance Sheet

December 31, 2018

2018

2017

Cash and cash equivalents

$55,000

$33,000

Net patient accounts receivable

20,000

30,000

Inventories

20,000

11,000

     Total current assets

$95,000

$74,000

Property and Equipment

290,000

70,000

Less: Accumulated depreciation

(14,000)

(6,000)

Net property and equipment

276,000

64,000

Long-term investments

27,000

0

Total assets

$398,000

$138,000

Accounts payable

$28,000

$12,000

Accrued expenses

6,000

8,000

   Total current liabilities

34,000

20,000

Long-term debt

130,000

20,000

Total Liabilities

$164,000

$40,000

Net Assets

234,000

98,000

Total liabilities and equity

$398,000

$138,000

Additional information: The organization did not sell any fixed assets in 2018

a. Prepare a properly formatted statement of cash flows for 2018 using the indirect method. See Exhibit 4.6.

b. How does Tampa Clinic’s statement of cash flows compare with Sunnyvale’s statement? See textbook, Exhibit 4.6.
Answer #1
a) STATEMENT OF CASH FLOWS
Cash flows from operating activities:
Net income 136000
Adjustments to reconcile net income to net cash profived by operations:
Depreciation expense 8000
Decrease in accounts receivable = 30000-20000 = 10000
Increase in inventory = 20000-11000 = -9000
Decrease in income accruals = 8000-6000 = -2000
Increase in accounts payable = 28000-12000 = 16000 23000
Net cash provided by operating activities 159000
Cash flows from investing activities:
Purchase of long term investments -27000
Purchase of plant and equipment (290000-70000) -220000
Net cash used in investing activities -247000
Cash flows from financing activities:
Cash from long term debt (130000-20000) 110000
Net cash used in financing activities 110000
Net increase (decrease) in cash 22000
Cash balance at beginning of year 33000
Cash balance at end of year 55000
b) Exhibit of Sunnvale not uploaded.